OpenStudy (anonymous):

The sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second. How many ergs of radiant energy does the sun produce in 1.55 ⋅ 10^7 seconds? (5 points)

OpenStudy (anonymous):

Will medal and fan

OpenStudy (anonymous):

ok lets take this problem step by step first lets think of 3.9*10^33 like one number just to make things easier this number is only written this way because the number is so large that the decimal point in 3.9 would have to move 33 places to the left to get the number the problem states "sun produces 3.9 ⋅ 10^33 ergs of radiant energy per second" now lets put it into a simple equation IMPORTANT MATH FACT whenever you see PER in a math problem it means multiplication (quick example: 3 apples PER person. if you have 3 people it would be 3*(the per)3 so 3.9*10^33 * 1second = amount of ergs just like the problem dictates now the problem says "How many ergs of radiant energy does the sun produce in 1.55 ⋅ 10^7 seconds" lets simplify this complex statement to how much does the sun produce in 1.55*10^7 seconds again think of 1.55*10^7 as one number now lets go back to our equation :3.9*10^33*1second = amount of ergs and replace 1 second with the amount of seconds the problem tells us to use 1.55*10^7 this brings us to the equation : 3.9*10^33* 1.55*10^7=amount of ergs you can do the rest!

OpenStudy (anonymous):

sorry just ignore the question marks in my reply idek why they appear